Circular No. 8235

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003jg IN NGC 2997 R. Martin, Perth Observatory, reports the discovery, on CCD images

taken on Oct. 24.800 and 28.752 UT with the 0.61-m Perth/Lowell Automated Telescope in the course of the Perth Automated Supernova Search, of an apparent supernova (red mag 17) located in NGC 2997. J. Biggs, Perth Observatory, writes that three unfiltered exposures taken with the 25-cm Mike Candy Telescope around Oct. 29.778 show the new object (R = 17 ± 1) at the following position: α = 9^{h}45^{m}37^{s}.91 ± 0′′.3, δ = −31^{o}11′21′′.0 ± 1′′.0 (equinox 2000.0), which is ≈ 11′′ east and 6′′ north of the galactic nucleus. Martin adds that nothing is visible at this location on an image taken on Sept. 25.876 (limiting mag 19), but the object does appear, just above the noise threshold, in an image taken on Oct. 5.850.

2003 QY_{90}

J. L. Elliot, Massachusetts Institute of Technology (MIT), reports that Sloan-r′-band observations, obtained by S. D. Kern (MIT) at the 6.5-m Clay telescope (+ MagIC) on Oct. 23 UT in ∼ 0′′.5 seeing, reveal highly elongated images of 2003 QY_{90} (cf. MPEC 2003-Q58). The consistency of the images over the 1.9-hr interval between the first and last frames rules out a blend with a background field object. Interpreting each elongated image as a pair of unresolved sources, analysis by K. B. Clancy (MIT), Kern, and Elliot implies the following properties: both objects are nearly identical in brightness (Δr′ = 0.1 ± 0.2); the separation of the pair is 0′′.34 ± 0′′.02; and the line between the two putative bodies has an approximate position angle of 129^{o} ± 5^{o}. Additional photometric and astrometric observations of 2003 QY_{90} are strongly encouraged.

V838 MONOCEROTIS M. Tapia, Instituto de Astronomía, Universidad Nacional Autonoma

de México; and P. Persi, Istituto Astrofisica Spaziale e Fisica Cosmica, Rome, report the following magnitudes observed with the 2.1-m telescope (+ mid-infrared camera CID) at San Pedro Martir on Oct. 5.5 UT: [8.7 μm] = 0.20, [9.7 μm] = −0.22, [12.5 μm] = −0.83, [19 μm] = −1.37. The object is cooler and has brightened at these wavelengths by ∼ 2 magnitudes since 2002 Sept. (cf. IAUC 7976) when observed with the same instrument.

2003 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green